OpenStudy (anonymous):

Concrete can be purchased by the cubic yard. How much will it cost to pour a slab 17 feet by 17 feet by 2 inches for a patio if the concrete costs $40.00 per cubic yard? a. $76.45 b. $71.36 c. $90.67 d. $87.56

OpenStudy (anonymous):

@_greatmath7

OpenStudy (anonymous):

B. First convert the patio slab into feet. We know that there are 12 inches in 1 feet. We convert 2inches into units of feet 17ft X 17ft X 2 in (1ft/12inches) = 48.2 cubic feet We then convert 48.2 ft^3 into cubic yard. We know there is 3 feet in 1 yard. 48.2 ft^3 X ( 1 yard / 3 feet)^3 = 48.2 ft^3 X (1 yard^3 / 27 ft^3) = 1.79 cubic yard To figure out the cost. We have 1.79 cubic yard for the patio. 1.79 yard^3 X ($40/yard^3) = $71.4

OpenStudy (anonymous):

I just rounded

OpenStudy (anonymous):

wow thank you so much!!!! you explained it so well.

OpenStudy (anonymous):

thanks and you are so welcome :)

OpenStudy (anonymous):

just tag me if you need anymore help! ;)

OpenStudy (mathstudent55):

You need to find how many cubic yards is the volume of the slab. Then multiply the cost per cubic yard by the number of cubic yards. \(V = LWH = \left( 17~ft \times \dfrac{yd}{3~ft} \right) \times \left(17~ft \times \dfrac{yd}{3~ft} \right) \times \left( 2~in. \times \dfrac{yd}{36~in.} \right) = 1.78 ~yd^3\) \(Cost = V \times (cost ~per ~yd^3) = 1.78 ~yd^3 \times 40 \dfrac{$}{yd^3} = $71.36\)